OpenStudy (anonymous):

Given the expression: 1990 x 1991 - 1989 x 1990. What counting number is equivalent to the expression?

OpenStudy (anonymous):

1990 x 1991 - 1989 x 1990 = 1990 ( 1991-1989) = 3980

OpenStudy (anonymous):

Thanks ... got it... y cant I think like you guys...!